In a survey of 320 doctors. 50% more favored Brand X pain reliever than Brand y. How many of the doctors favored Brand X?
Answer by stanbon(75887)

Let number that favor brand y be "a"
Then number that favor brand x will be "a + 50%a" = 1.5a
EQUATION:
number favor y + number favor x = 320
a + 1.5a = 320
10a + 15a = 3200
25a = 3200
a = 128 (number who favor y)
1.5a = 192 (number who favor x)
Cheers,
Stan H.